OpenStudy (hdrager):

y=5/8sin(-8pi/3 x) Find the period

OpenStudy (anonymous):

The period is equal to 2pi/B in the equation Asin(B(x+C))+D, so in this equation B is -8pi/3. So the period is equal to (2pi)/(-8pi/3) which is -6/8.

OpenStudy (hdrager):

@trainwrecking could you show me the process of getting the answer?

Directrix (directrix):

Study Chart

OpenStudy (hdrager):

@Directrix I know how to do the problem, I just didn't get the correct answer. I want to know where exactly I went wrong if possible

Directrix (directrix):

@trainwrecking Is the period + 3/4 rather than -3/4?

OpenStudy (anonymous):

@Directrix You're totally right, I left out the absolute value, thanks! @hdrager \[\frac{ 2\pi }{ \left| \frac{ -8\pi }{ 3 } \right|} = 2\pi*\frac{ 3 }{ 8\pi } = \frac{ 6 }{ 8 } = \frac{ 3 }{ 4 }\] Does that help?

OpenStudy (hdrager):

@trainwrecking Yes thank you I get where I went wrong now

Directrix (directrix):

To get the period: 1) Write the equation in this form: F(x) = A*sin(B*x – C) + D 2) Period of the sine function will be the "regular period" of the sine function ( 2 pi) divided by the absolute value of the coefficient of the x term, B.